I would like to post the limit of 4 questions. I am struggling with the class and could use all the help I can get.
1. If you are in a rocket going perfectly vertical into the air what is the slope that you are encountering
2. The line parallel to the lin y=3x+5 has what slope
3. The line perpendivular to the line y=-x/3 + 5 has what slope
4. write the equation for the line y=-x/3+5 in standard form

Answer by mukhopadhyay(490)

1. If you are in a rocket going perfectly vertical into the air what is the slope that you are encountering
Answer: Slope is 0; Since it is going vertical, the graph is parallel to the y-axis. The equation formed by this movement is y=k;
The co-efficient of x=0 in the above equation is 0=>slope=0
2. The line parallel to the lin y=3x+5 has what slope
Answer: Slope=3 (in y=mx+c, the slope is m)
3. The line perpendivular to the line y=-x/3 + 5 has what slope
Slope of the above line is -1/3, Slope of the perpendicular line is negative of the reciprocal of this, which is -1/(-1/3)= 3
4. write the equation for the line y=-x/3+5 in standard form
y=-x/3+5=>y=-1/3(x-15)=>3y=15-x=>x+3y-15=0
Answer:x+3y-15=0